# The impact of non-Gaussianity when searching for Primordial Black Holes with LISA

### v2

@@ 1 Introduction
Gravitational Waves (GWs) provide a unique window into both the early and late Universe. GW signals, in fact, are expected to be generated by a plenty of different astrophysical and cosmological processes (see [1, 2, 3] for some reviews), spanning a much wider frequency range, from the extremely high frequencies probed indirectly by Cosmic Microwave Background (CMB) experiments ($\sim 10^{-18}$ Hz) up to the GHz regime [4, 5, 6, 7, 8]. To access this broader spectrum, several detectors have been proposed. Among them, ground-based interferometers like the Einstein Telescope (ET) [9] and Cosmic Explorer (CE) [10] aim to improve sensitivity down to $\sim 0.1$ Hz, while the space-based Laser Interferometer Space Antenna (LISA) [11] will target the milli-Hz band. LISA, who is expected to be operative within a decade, will feature a triangular configuration and will probe frequencies that bridge the gap between ground-based interferometers and those accessible via Pulsar Timing Arrays (PTA) or CMB experiments. Importantly, many cosmological models predict a measurable gravitational wave background in this intermediate frequency band, see e.g. [12, 3, 13, 1]. As a result, any detection or non-detection by LISA will carry fundamental implications for the understanding of early-universe processes and fundamental physics [4, 14, 15].
@@
Among other sources, scalar-induced gravitational waves (SIGW) [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] constitute a particularly interesting cosmological GW background detectable by LISA (see also [27] for a recent physical explanation of the origin of the signal). As the name suggests, these signals are generated from scalar fluctuations, which act as sources for tensor modes at second order in perturbation theory. As the shape and amplitude of the resulting SIGW spectrum are derived from the scalar power spectrum [28], its detection would be fundamental to unveil the dynamics of the early-Universe on scales much smaller than the ones probed by Planck and more recently by the ACT and SPT collaborations [29, 30, 31]. Various inflationary models [32, 33, 34] [^1] Examples of this are single-field scenarios where the potential shows specific features which lead to an ultra-slow-roll (USR) phase [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] as well as multi-field models [32, 33, 46, 47, 48, 49, 50, 51, 52, 53], but also mechanisms such as preheating or early matter-dominated eras which can produce a strong enhancement of the GW spectrum without the need of a largely enhance of the scalar power spectrum [54, 55, 56, 57, 58]., in fact, predict an amplification of the scalar fluctuations around LISA scales with respect to the nearly-flat power spectrum on CMB scales [59] and this enhancement could possibly generate a large amount of SIGW. In addition, SIGWs are directly sourced by the four-point function of the scalar fluctuation and are sensitive to the possible presence of non-Gaussianity (NG) in the scalar perturbation, as it has been recently analysed by [60, 61, 23, 24, 62]. Primordial NG represents a further powerful probe of the early Universe, being sensitive to the dynamics and interactions in the early Universe. Although standard inflationary scenarios predict mild deviations from Gaussian behaviour, other models could generate substantial NG perturbations [63, 64, 65]; thus, constraining primordial NG is fundamental for discriminating between models. Nevertheless, Planck has been able to put tight constraints on the amount of NG, typically parametrized by the quantity $f_{\rm NL}$, on large scales [66], but small scales remain unconstrained.
@@
Interestingly, the same perturbations that generate SIGW are responsible for the formation of primordial BH (PBH) [67, 68, 69, 70, 71]. Different mechanisms have been proposed to possibly produce PBHs, but in the standard formation scenario, which we consider in this work, PBHs are formed as a consequence of the gravitational collapse of large overdensities after they re-enter the horizon. In this scenario, the formation of PBHs is boosted if the scalar power spectrum is enhanced, resulting in a non-negligible PBH abundance [35, 72, 73]. The mass of the resulting GWs is linked to the wavenumber associated with the perturbation responsible for its formation. Thus, to SIGWs in the millihertz frequency band correspond PBHs in the asteroidal-mass window, which is, as of today, fully unbounded [74, 75, 76, 77, 78]. Therefore, detecting or setting upper bounds on SIGWs would have fundamental implications for constraining the PBH abundance.
@@
Previous works have explored how a possible PTA observation of SIGWs would impact constraints on the PBH abundance (see, e.g., [79, 80]). In this work, our aim is to understand the implications of SIGW in the LISA frequency range on the abundance of PBHs in the asteroidal mass range. By analysing the possible detection of SIGW over the whole range, we derive the corresponding bounds on the amplitude of the scalar power spectrum, first assuming Gaussianity of the primordial perturbations and then considering the presence of NG. Thus, we derive the bounds on the PBH abundance by studying the effect of different widths of the primordial power spectrum and the presence of astrophysical foregrounds in the LISA range. After that, we discuss the detectability of primordial NG and the mutual implications between SIGWs and PBH. The paper is organised as follows. In Section 2, we introduce primordial NG, focussing on the local shape, and explore its impact on the computation of SIGWs and PBH abundance, including a brief discussion of the perturbativity criterion. In Section 3, we describe how to relate the projected constraints on SIGWs to the PBH abundance. We then present our results in Section 4, where we also discuss the uncertainties in LISA’s capability to constrain primordial NG. We conclude in Section 5.
@@ 2 Non-Gaussian primordial curvature perturbations
where the term $\langle\zeta_{G}^{2}\rangle$ has been introduced in order to let $\langle\zeta\rangle=0$. The amount of NG can thus be typically quantified by $f_{\rm NL}$. Its value depends on the specific inflation model considered. Although the Planck Collaboration [66] has placed strong constraints on different types of NG, all of which are still compatible with 0, the error bars are still large. We focus in this work on NG of the local type, which is typical of multi-field inflation models (see, e.g. [81]); moreover, its imprints on the SIGWs have been widely studied, given the feasibility of the calculations [23, 24].
@@ 2.1 Scalar-induced gravitational waves
where the linear evolution is contained in the integration kernel $I$, whose shape depends on the expansion history of the Universe [87, 88]. We will assume radiation domination for the rest of this paper. However, modifications to it relevant in the LISA range are, for instance, possible due to brief early matter-dominated periods [89, 90, 91, 92, 93, 94, 95, 96, 97] or the softening of the equations of state due to a first-order electroweak phase transition [98, 99, 100, 101]. The SIGW energy density per logarithmic $k$ at time $\eta$ is
@@ 2.2 Perturbativity
Figure 1: Comparison between the perturbative expansion of the non-Gaussian field $\zeta$ as in Eq. (2.2) and the full functional form $\zeta=F(\zeta_{G})$ for two benchmark scenarios: USR and curvaton models. The coefficients $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are shown for $f_{\mathrm{NL}}=1$ (top) and $f_{\mathrm{NL}}=10$ (bottom).
@@
$\mathcal{C}_{s}=\frac{1}{s!(2A)^{\frac{s+1}{2}}}\int\frac{{\rm d}\zeta_{G}}{\sqrt{\pi}}e^{-\zeta_{G}^{2}/(2A)}\,H_{n}\left(\frac{\zeta_{G}}{\sqrt{2A}}\right)F(\zeta_{G})\,$, (2.17)
effectively perform a resummation over Wick contractions at a single point [103], $H_{n}$ are Hermite polynomials and
@@
is the variance of fluctuations at a point in space.
@@
with $F_{\rm NL,1}=1$ and $F_{\rm NL,2}=(3/5)f_{\rm NL}$, etc. The coefficients (2.17) can then be expressed by the series by analytically evaluating the integrals (see [106] for an alternative derivation)
@@
where $\beta$ is related to inflationary dynamics. For the curvaton, the function $F$ is significantly more complicated (see, e.g., section 2.4 of [50]). Nevertheless, the non-Gaussianity is controlled by the curvaton decay parameter $r_{\rm dec}$ also in that case. To relate these examples with the perturbative Eq. (2.2), we will match the quadratic terms, which gives, e.g, that $f_{\rm NL}=(5/6)\beta$ in the USR.
@@
The radius of convergence of this series is 0, i.e., it diverges. Thus, to obtain a comparison with the non-perturbative expression, we must compute the integral in Eq. (2.17) numerically. Nevertheless, the appearance of divergent perturbative series is commonplace in field theory and (2.23) should be interpreted as an asymptotic series whose breakdown can be estimated by when the terms begin to grow. By using the Stirling approximation, we find that terms in the series (2.23) begin to grow when $n\gtrsim 1/(2A\beta^{2})e^{-(s-1)/n}$, thus demanding that at least the first 3 terms in the series decrease at the order $f_{\rm NL}$ ($s=2$), we find that $A\beta^{2}\lesssim 0.12$ which translates to
@@
The estimate perturbativity is somewhat arbitrary. We choose a criterion that quantifies how accurately the perturbative estimate $C_{s}^{\rm p}$ in Eq. (2.21) matches the non-perturbative resummed estimate $C_{s}^{\rm F}$ in Eq. (2.17). For concreteness, we characterize perturbativity by
$2/3<C_{s}^{\rm p}/C_{s}^{\rm F}<3/2$ (2.25)
and the breakdown of perturbativity by the violation of this condition. The perturbativity bounds are shown in Fig. 2. We find that, for the USR case, it breaks down for $Af^{2}_{\rm NL}\gtrsim 0.06$, which is in good agreement with Eq. (2.15), while for the curvaton case, it depends on $f_{\rm NL}$.
Indeed, for $f_{\rm NL}=1$, the curvaton model becomes non-perturbative at values comparable to USR, namely when $Af_{\rm NL}^{2}\gtrsim 0.06$. For larger $f_{\rm NL}\gg 1$, this bound is much weaker, requiring only $Af_{\rm NL}^{2}\gtrsim 1$. Curiously, we also find that the perturbativity criterion becomes much stronger when $|f_{\rm NL}|\ll 1$. Although this result seems counterintuitive, it is possible to understand it by considering the intricate shape of $F$ for the curvaton. Importantly, the curvaton does not possess a Gaussian limit and the cubic and higher order terms do not vanish when $f_{\rm NL}\to 0$. This can already be observed when considering the next-to-leading order, that is $g_{\rm NL}$. In terms of the curvaton decay parameter $r_{\rm dec}$, we have that [109]
@@ 2.3 Primordial black hole abundance
where
@@
Figure 3: Constraints and prospective sensitives on the primordial scalar power spectrum for current and future GW experiments. The curvature perturbations are assumed Gaussian statistics and to follow log-normal power spectra with for two different widths, $\Delta=0.$ (solid lines) and $\Delta=0.9$ (dashed lines). The existing constraints for Planck+ACT[127] and Ly -$\alpha$[128] are taken directly from their respective analyses, while for FIRAS[129, 130, 80] and NANOGrav15[131, 132, 80] they are computed using the latest data and assuming our template for the primordial power spectrum. The sensitivities are shown for SKA [133], LISA [11], AEDGE [134], TianQin [135], ET [9], CE [10] and aLIGO [136] estimated using ${\rm SNR}_{\rm th}=10$ and an observation time of 11 year. The light grey lines indicate the upper bound on AA due to the PBH overproduction (assuming the TS formalism).
@@ 3.1 Constraints on the scalar power spectrum
Fig. 3 shows the prospective sensitivities for Gaussian curvature perturbations assuming an observation time $T_{\rm obs}=1$ [^5] Prospects for a different SNR and $T_{\rm obs}$ can be obtained by multiplying the bounds on $A$ by $({\rm SNR}^{\rm new}/{{\rm SNR}^{\rm old}})^{1/2}(T^{\rm new}_{\rm obs}/T^{\rm old}_{\rm obs})^{-1/4}$. together with current constraints on the amplitude of the primordial power spectrum. At the largest scales, we report the constraints from Planck+ACT [127] and Lyman-$\alpha$ (Ly-$\alpha$[137] which do not depend on our templates for the power spectrum. For higher $k$, the constraints and sensitivities are evaluated for two different log-normal curvature power spectra with a narrow ($\Delta=0.1$) and a wide peak ($\Delta=0.9$). The constraints from Firas [129, 130] and NANOGrav [131, 132] are obtained as in [80]. Finally, we show the projected sensitivities of SKA [133], LISA [11], AEDGE [134], TIANQin [135], ET [9], CE [10] and aLIGO [136]. For the prospects from PTA [138] and SKA [139] we neglected astrophysical foregrounds due to SMBH binaries.

### v3

@@ 1 Introduction
Gravitational Waves (GWs) provide a unique window into both the early and late Universe. GW signals, in fact, are expected to be generated by a variety of different astrophysical and cosmological processes (see [52, 156, 99] for some reviews), spanning a much wider frequency range, from the extremely low frequencies probed indirectly by Cosmic Microwave Background (CMB) experiments ($\sim 10^{-18}$ Hz) up to the GHz regime [20, 43, 1, 8, 9]. To access this broader spectrum, several detectors have been proposed. Among them, ground-based interferometers like the Einstein Telescope (ET) [127] and Cosmic Explorer (CE) [157] aim to improve sensitivity down to $\sim 0.1$ Hz, while the space-based Laser Interferometer Space Antenna (LISA) [15] will target the milli-Hz band. LISA, which is expected to be operational within a decade, will feature a triangular configuration and will probe frequencies that bridge the gap between ground-based interferometers and those accessible via Pulsar Timing Arrays (PTA) or CMB experiments. Importantly, many cosmological models predict a measurable gravitational wave background in this intermediate frequency band, see e.g. [128, 99, 45, 52]. As a result, any detection or non-detection by LISA will carry fundamental implications for the understanding of early-universe processes and fundamental physics [20, 18, 162].
@@
Among other sources, scalar-induced gravitational waves (SIGW) [130, 31, 131, 54, 73, 77, 16, 5, 148, 44, 4] constitute a particularly interesting cosmological GW background detectable by LISA (see also [104] for a recent physical explanation of the origin of the signal). As the name suggests, these signals are generated from scalar fluctuations, which act as sources for tensor modes at second order in perturbation theory. As the shape and amplitude of the resulting SIGW spectrum are derived from the scalar power spectrum [149], its detection would be fundamental to unveiling the dynamics of the early-Universe on scales much smaller than those probed by Planck and more recently by the ACT and SPT collaborations [10, 125, 50]. Various inflationary models [95, 112, 145] [^1] Examples of this are single-field scenarios where the potential shows specific features that lead to an ultra-slow-roll (USR) phase [106, 115, 102, 116, 129, 135, 143, 111, 25, 13, 14] as well as multi-field models [95, 112, 151, 110, 41, 164, 88, 165, 173, 101], and also mechanisms such as preheating or early matter-dominated eras that can produce a strong enhancement of the GW spectrum without the need for a large enhancement of the scalar power spectrum [109, 75, 146, 147, 27]. in fact, they predict an amplification of the scalar fluctuations around LISA scales with respect to the nearly-flat power spectrum on CMB scales [12], and this enhancement could possibly generate a large amount of SIGW. In addition, SIGWs are directly sourced by the four-point function of the scalar fluctuation and are sensitive to the possible presence of non-Gaussianity (NG) in the scalar perturbation, as it has been recently analysed by [46, 169, 176, 5, 148, 179]. Primordial NG represents a further powerful probe of the early Universe, being sensitive to the dynamics and interactions in the early Universe. Although standard inflationary scenarios predict mild deviations from Gaussian behaviour, other models could generate substantial NG perturbations [118, 23, 30]; thus, constraining primordial NG is fundamental for discriminating between models. Nevertheless, Planck has been able to put tight constraints on the amount of NG, typically parametrized by the quantity $f_{\rm NL}$, on large scales [11], but small scales remain unconstrained.
@@
Interestingly, the same perturbations that generate SIGW are responsible for the formation of primordial BH (PBH) [178, 100, 56, 58, 63]. Different mechanisms have been proposed to possibly produce PBHs, but in the standard formation scenario, which we consider in this work, PBHs are formed as a consequence of the gravitational collapse of large overdensities after they re-enter the horizon. In this scenario, the formation of PBHs is boosted if the scalar power spectrum is enhanced, resulting in a non-negligible PBH abundance [106, 107, 38]. The mass of the resulting GWs is linked to the wavenumber associated with the perturbation responsible for its formation. Thus, to SIGWs in the mHz frequency band correspond PBHs in the asteroidal-mass window, which is, as of today, fully unbounded [28, 29, 175, 24, 26]. Therefore, detecting or setting upper bounds on SIGWs would have fundamental implications for constraining the PBH abundance.
@@
Previous works have explored how a possible PTA observation of SIGWs would impact constraints on the PBH abundance (see, e.g., [64, 105]). In this work, our aim is to understand the implications of SIGW in the LISA frequency range on the abundance of PBHs in the asteroidal mass range. By analysing the possible detection of SIGW over the whole range, we derive the corresponding bounds on the amplitude of the scalar power spectrum, first assuming Gaussianity of the primordial perturbations and then considering the presence of NG. Thus, we derive the bounds on the PBH abundance by studying the effect of different widths of the primordial power spectrum and the presence of astrophysical foregrounds in the LISA range. After that, we discuss the detectability of primordial NG and the mutual implications between SIGWs and PBH. The paper is organised as follows. In Section 2, we introduce primordial NG, focusing on the local shape, and explore its impact on the computation of SIGWs and PBH abundance, including a brief discussion of the perturbativity criterion. In Section 3, we describe how to relate the projected constraints on SIGWs to the PBH abundance. We then present our results in Section 4, where we also discuss the uncertainties in LISA’s capability to constrain primordial NG. We conclude in Section 5.
@@ 2 Non-Gaussian primordial curvature perturbations
where the term $\langle\zeta_{G}^{2}\rangle$ has been introduced in order to let $\langle\zeta\rangle=0$. The amount of NG can thus be typically quantified by $f_{\rm NL}$. Its value depends on the specific inflation model considered. Although the Planck Collaboration [11] has placed strong constraints on different types of NG, all of which are still compatible with 0, the error bars are still large. We focus in this work on NG of the local type, which is typical of multi-field inflation models (see, e.g. [171]); moreover, its imprints on the SIGWs have been widely studied, given the feasibility of the calculations [46, 5, 148].
@@ 2.1 Scalar-induced gravitational waves
where the linear evolution is contained in the integration kernel $I$, whose shape depends on the expansion history of the Universe [117, 76]. We will assume radiation domination for the rest of this paper. However, modifications to it relevant in the LISA range are, for instance, possible due to brief early matter-dominated periods [90, 144, 155, 47, 69, 96, 62, 48, 74] or the softening of the equations of state due to a first-order electroweak phase transition [53, 83, 81, 122]. The SIGW energy density per logarithmic $k$ at time $\eta$ is [^2] Note that in our convention the tensor perturbation at second order is defined as $h_{ij}/2$.
@@ 2.2 Perturbativity
Figure 1: Ratio of the coefficients $\mathcal{C}_{s}$ obtained with the perturbative expansion of the non-Gaussian field $\zeta$ as in Eq. (2.2) and the full functional form $\zeta=F(\zeta_{G})$ for two benchmark scenarios: USR (solid lines) and curvaton (dashed lines) models. The colors, red and blue, indicate the order of the coefficient chosen, respectively $s=1$ and $s=2$. The coefficients are shown for $f_{\mathrm{NL}}=1$ (left) and $f_{\mathrm{NL}}=10$ (right).
@@
$\mathcal{C}_{s}=\frac{1}{s!(2A)^{\frac{s+1}{2}}}\int\frac{{\rm d}\zeta_{G}}{\sqrt{\pi}}e^{-\zeta_{G}^{2}/(2A)}\,H_{s}\left(\frac{\zeta_{G}}{\sqrt{2A}}\right)F(\zeta_{G})\,$, (2.17)
effectively perform a resummation over Wick contractions at a single point, $H_{n}$ are Hermite polynomials and
@@
is the variance of fluctuations at a point in space. This non-perturbative approach relies on the conservation of probability when mapping $\zeta$ to $\zeta_{G}$ and the decomposition of the joint Gaussian probability distribution into Hermite polynomials which allows one to write any $n$-point correlator of $\zeta$ entirely in terms of the coefficients $\mathcal{C}_{s}$ and combinations of the two-point function of the Gaussian field (for details, see [170]). Importantly, as (2.17) does not rely on the series expansion of $F(\zeta_{G})$, it can be applied even when $F(\zeta_{G})$ is not analytic.
@@
with $F_{\rm NL,1}=1$ and $F_{\rm NL,2}=(3/5)f_{\rm NL}$, etc. The coefficients (2.17) can then be expressed by the series by evaluating the integrals (see [124] for an alternative derivation)
@@
where $\beta$ is related to the subsequent constant-roll phase determined by the inflationary dynamics [^4] For simplicity, in USR scenarios we neglect the contribution to the PBHs abundance coming from the type II subdominant channel [79, 172, 80].. For the curvaton, the function $F$ is significantly more complicated (see, e.g., section 2.4 of [88]). Nevertheless, the non-Gaussianity is controlled by the curvaton decay parameter $r_{\rm dec}$ also in that case. To relate these examples with the perturbative Eq. (2.2), we will match the quadratic terms, which gives, e.g, that $f_{\rm NL}=(5/6)\beta$ in the USR.
@@
The radius of convergence of this series is 0, i.e., it diverges. Thus, to obtain a comparison with the non-perturbative expression, we must compute the integral in Eq. (2.17) numerically. Nevertheless, the appearance of divergent perturbative series is commonplace in field theory and (2.23) should be interpreted as an asymptotic series whose breakdown can be estimated by when the terms begin to grow. By using the Stirling approximation, we find that terms in the series (2.23) begin to grow when $n\gtrsim 1/(2A\beta^{2})e^{-(s-1)/n}$, thus demanding that at least the first 3 terms in the series decrease at the order $f_{\rm NL}$ ($s=2$), we find that $A\beta^{2}\lesssim 0.12$, which translates to
@@
The estimate of perturbativity is somewhat arbitrary. We choose a criterion that quantifies how accurately the perturbative estimate $\mathcal{C}_{s}^{p}$ in Eq. (2.21) matches the non-perturbative resummed estimate $\mathcal{C}_{s}^{F}$ in Eq. (2.17). For concreteness, we characterize perturbativity by demanding that
$2/3<\mathcal{C}_{s}^{p}/\mathcal{C}_{s}^{F}<3/2,$ (2.25)
and the breakdown of perturbativity is indicated by the violation of this condition. The perturbativity bounds are shown in Fig. 2. We find that, for the USR case, it breaks down for $Af^{2}_{\rm NL}\gtrsim 0.06$, which is in good agreement with Eq. (2.15), while for the curvaton case, it depends on $f_{\rm NL}$.
Indeed, for $f_{\rm NL}=1$, the curvaton model becomes non-perturbative at values comparable to USR, namely when $Af_{\rm NL}^{2}\gtrsim 0.06$. For larger $f_{\rm NL}\gg 1$, this bound is much weaker, requiring only $Af_{\rm NL}^{2}\gtrsim 1$. Curiously, we also find that the perturbativity criterion becomes much stronger when $|f_{\rm NL}|\ll 1$. Although this result seems counterintuitive, it is possible to understand it by considering the intricate shape of $F$ for the curvaton. Importantly, the curvaton does not possess a Gaussian limit, and the cubic and higher order terms do not vanish when $f_{\rm NL}\to 0$. This can already be observed when considering the next-to-leading order, that is, $g_{\rm NL}$. In terms of the curvaton decay parameter $r_{\rm dec}$, we have that [160]
@@ 2.3 Primordial black hole abundance
where $\mathcal{C}$ is the so called compaction function, generically defined as twice the local mass excess over the areal radius, and
@@
Figure 3: Constraints and prospective sensitives on the primordial scalar power spectrum for current and future GW experiments. The curvature perturbations are assumed Gaussian statistics and to follow log-normal power spectra with for two different widths, $\Delta=0.1$ (solid lines) and $\Delta=0.9$ (dashed lines). The existing constraints for Planck+ACT[49] and Ly -$\alpha$[35] are taken directly from their respective analyses, while for FIRAS[65, 66, 105] and NANOGrav15[6, 7, 105] they are computed using the latest data and assuming our template for the primordial power spectrum. The sensitivities are shown for SKA [108], LISA [15], AEDGE [78], TianQin [126], ET [127], CE [157] and aLIGO [3] estimated using ${\rm SNR}_{\rm th}=10$ and an observation time of 11 year. The light grey lines indicate the upper bound on AA due to the PBH overproduction (assuming the TS formalism).
@@ 3.1 Constraints on the scalar power spectrum
Fig. 3 shows the prospective sensitivities for Gaussian curvature perturbations assuming an observation time $T_{\rm obs}=1$ yr [^7] Prospects for a different SNR and $T_{\rm obs}$ can be obtained by multiplying the bounds on $A$ by $({\rm SNR}^{\rm new}/{{\rm SNR}^{\rm old}})^{1/2}(T^{\rm new}_{\rm obs}/T^{\rm old}_{\rm obs})^{-1/4}$. together with current constraints on the amplitude of the primordial power spectrum. At the largest scales, we report the constraints from Planck+ACT [49] and Lyman-$\alpha$ (Ly-$\alpha$[36] which do not depend on our templates for the power spectrum. For higher $k$, the constraints and sensitivities are evaluated for two different log-normal curvature power spectra with a narrow ($\Delta=0.1$) and a wide peak ($\Delta=0.9$). The constraints from Firas [65, 66] and NANOGrav [6, 7] are obtained as in [105]. Finally, we show the projected sensitivities of SKA [108], LISA [15], AEDGE [78], TIANQin [126], ET [127], CE [157] and aLIGO [3]. For the prospects from PTA [61] and SKA [22] we neglected astrophysical foregrounds due to SMBH binaries.